OpenStudy (credmond):

Hey guys! I just need help finishing a problem! y=e^x(secx) So far I have u= e^x. u'=e^x, v=secx, v'=secxtanx So I set it up as (e^x)(secxtanx)+(e^x)(secx) Need help simplifying!

ganeshie8 (ganeshie8):

are you trying to find the derivative of y ?

OpenStudy (credmond):

ohhh sorry yes!

ganeshie8 (ganeshie8):

looks you have applied the product rule correctly

ganeshie8 (ganeshie8):

you're done with calc part

OpenStudy (credmond):

Right! I just don't know how to simplify this :/

OpenStudy (anonymous):

you can take \(e^x \cdot \sec(x)\) common for that.. :)

OpenStudy (anonymous):

*from, not for.. :)

OpenStudy (anonymous):

But this will not be called as Simplifying it, it will be just rearrangement only.. :P

OpenStudy (credmond):

so I guess I'm just wondering what the final product be?

ganeshie8 (ganeshie8):

putting the final answer in a nice format is good, but is ur prof particular about niceness ?

OpenStudy (anonymous):

I think we can just rearrange it only,,

ganeshie8 (ganeshie8):

if you're just asked to find the derivative, you're done right after applying the product rule

OpenStudy (credmond):

Ohhh okay. On my test, he said I was okay until that point. I tried to simplify it farther and he marked points off. So maybe that's what he meant!? To just stop after the product rule?

ganeshie8 (ganeshie8):

lol he should not cut marks just because you made it look better

ganeshie8 (ganeshie8):

anyways he is correct in someways, this is calc course, focus should be on learning calc concepts...

OpenStudy (credmond):

Haha true true. While I got you here, do you mind checking to see if I took this other problem as long as I could, as well?

ganeshie8 (ganeshie8):

sure, ask wil try..

OpenStudy (credmond):

y=tanx/x u=tanx u'=sec^2(x) v=x v'=1 So I put it all together and got (x) (sec^2(x) - (1) (tanx) / x^2

ganeshie8 (ganeshie8):

looks good, but nobody leaves (1)(tanx) like that :o

OpenStudy (credmond):

oh. right

OpenStudy (credmond):

soooo (x)(sec^2(x)) - (tanx) / x^2

OpenStudy (credmond):

It can't be simplified any further right?

ganeshie8 (ganeshie8):

(xsec^2x - tanx) / x^2 looks good as final answer

ganeshie8 (ganeshie8):

note that parenthesis should be there for the entire numerator

OpenStudy (credmond):

Gotcha! Great!!! Thank you so much! :)

ganeshie8 (ganeshie8):

np :)